OpenStudy (anonymous):

An angle measure of 234 degrees is equivalent to ____ radians. Round your answer to the nearest hundredth when necessary

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large 234~~degrees \) \(\large 234~~degrees \times \frac{\pi~~radians}{180~~degrees}\) \(\large 234 \times \frac{\pi~~radians}{180}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

I got 4.082

ganeshie8 (ganeshie8):

correct ! round it to hundredth

ganeshie8 (ganeshie8):

4.082 ^

OpenStudy (anonymous):

thanks!! :)

ganeshie8 (ganeshie8):

when u round it to hundredth, it becomes 4.08